OpenStudy (anonymous):

If 12% of an iceberg is above water, what fraction is in the water?

OpenStudy (anonymous):

12% is 12/100, so the rest of the iceberg, 100-12= 88 is undewr water

OpenStudy (anonymous):

this is a trick question the whole thing is in the water! lololol

OpenStudy (anonymous):

So 88/100, or 88% is under water

OpenStudy (anonymous):

It's a multiple choice, and these were the answers that were given; 3/25 12/88 22/25 22/3

OpenStudy (anonymous):

22/25

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

divide both 88 and 100 by 4

OpenStudy (anonymous):

Thank you both very much <3

OpenStudy (anonymous):

It is the simplified version of 88/100

OpenStudy (anonymous):

Just happy to help :)

OpenStudy (anonymous):

Need anything else?

OpenStudy (anonymous):

I also have this one, if you wouldn't mind helping; In a basketball free-throw contest, four players take the same number of free throws. Player W makes 33 of 40 his free throws, Player X makes ¾ of his free throws, Player Y makes 85% of her free throws, and Player Z makes 4/5 of her free throws. Which player makes the greatest number of free throws?

OpenStudy (anonymous):

33/40=82.5%, 3/4 = 75%, 85%=85% (obviously :)) and 4/5= 80%. 85%>82.5%>80%>75%. Y made 85% of their shots, which is more than the others

OpenStudy (anonymous):

So Player Y made the greatest number of shots

OpenStudy (anonymous):

Thank you so much for all of your help. You're a lifesaver <3

OpenStudy (anonymous):

My pleasure. You can tag my name anytime you need help, and if I'm online, I'll do my best to help you :)